The sum of the measures of the interior angles of a polygon Is five times the sum of the measures of its exterior angles, one angle at each vertex. How many sides does the polygon have?
Answer by stanbon(75887)

The sum of the measures of the interior angles of a polygon Is five times the sum of the measures of its exterior angles, one angle at each vertex. How many sides does the polygon have?
---
The sum of the exterior angles is ALWAYS 360 degrees.
--
So, sum of the interior angles is 5*360 = 1800 degrees.
----
Let n be the number of sides:
Equation:
(n-2)180 = 1800
n-2 = 10
n = 12 sides
====================
Cheers,
Stan H.
====================